Given the values of ΔH∘rxn, ΔS∘rxn, and T, determine ΔSuniv. Part A ΔH∘rxn=+84 kJ , ΔSrxn=+141 J/K , T= 308 K Express your answer as an integer. Part B ΔH∘rxn=+84 kJ , ΔSrxn=+141 J/K , T= 754 K Express your answer as an integer. Part C ΔH∘rxn=+84 kJ , ΔSrxn=− 141 J/K , T= 308 K Express your answer as an integer. Part D ΔH∘rxn=− 84 kJ , ΔSrxn=+ 141 J/K , T= 403 K Express your answer as an integer. Part E Predict whether or not the reaction in part A is spontaneous. Part F Predict whether or not the reaction in part B is spontaneous.Part G Predict whether or not the reaction in part C is spontaneous. Part H Predict whether or not the reaction in part D is spontaneous.
Answer #1

A)

Given:

ΔH = 84 KJ

= 8.4*10^4 J

T = 308 K

use:

ΔS Surr = -ΔH/T

= -(8.4*10^4)/308

= -2.727*10^2 J/K

ΔSuniv = ΔS Sys + ΔS Surr

= 141 J/K + -2.727*10^2 J/K

= -1.317*10^2 J/K

Answer: -132 J/K

B)

Given:

ΔH = 84 KJ

= 8.4*10^4 J

T = 754 K

use:

ΔS Surr = -ΔH/T

= -(8.4*10^4)/754

= -1.114*10^2 J/K

ΔSuniv = ΔS Sys + ΔS Surr

= 141 J/K + -1.114*10^2 J/K

= 29.59 J/K

Answer: 29.6 J/K

C)

Given:

ΔH = 84 KJ

= 8.4*10^4 J

T = 308 K

use:

ΔS Surr = -ΔH/T

= -(8.4*10^4)/308

= -2.727*10^2 J/K

ΔSuniv = ΔS Sys + ΔS Surr

= -141 J/K + -2.727*10^2 J/K

= -4.137*10^2 J/K

Answer: -414 J/K

D)

Given:

ΔH = -84 KJ

= -8.4*10^4 J

T = 403 K

use:

ΔS Surr = -ΔH/T

= -(-8.4*10^4)/403

= 2.084*10^2 J/K

ΔSuniv = ΔS Sys + ΔS Surr

= 141 J/K + 2.084*10^2 J/K

= 349 J/K

Only 4 parts at a time please